OpenStudy (anonymous):

H4P3: I found Rth but Vth is wrong ? How to find Vth

OpenStudy (nurali):

\[V _{TH}=I _{0}*\left( 1+\frac{ Z }{ R _{1} } \right)*\left[ \frac{ 1 }{ R _{1} }+\frac{ 1 }{ R _{2} }+\frac{ Z}{ R _{1}*R _{2} } \right]^{-1}\]

OpenStudy (anonymous):

yeah thats correct, it works for both Rth and Vth... thanx :)

OpenStudy (anonymous):

I solved this using the following equation: \[V _{th}*I_o\]

OpenStudy (anonymous):

Thank you Nurali. It is right

OpenStudy (nurali):

My Pleasure